# Identification and Optimal Nonlinear Control of Turbojet Engine Using Koopman Eigenfunction Model

### v2

@@ Abstract
Gas turbine engines represent complex highly nonlinear dynamical systems. Deriving their physics-based models can be challenging as it requires performance characteristics, that are not always available, and one often has to make many simplifying assumptions. In this paper, the limitations of conventional experimental methods used to derive component-level and locally linear parameter-varying models are discussed and addressed by employing identification techniques based on data collected from standard engine operation under closed-loop control. The rotor dynamics were estimated using the sparse identification of nonlinear dynamics. Subsequently, the autonomous part of the dynamics was mapped into an optimally constructed Koopman eigenfunction space. The process included eigenvalue optimization using metaheuristic algorithms and temporal projection, followed by gradient-based eigenfunction identification. The resulting Koopman model was validated against an in-house reference component-level model. A globally optimal nonlinear feedback controller and a Kalman estimator were then designed in the eigenfunction space and compared to the classical and gain-scheduled proportional-integral controllers, as well as a proposed internal model control approach. The eigenmode structure allowed targeting individual modes during the optimization process, resulting in a better performance tuning. The results showed that the Koopman-based controller outperformed the other benchmark controllers in both reference tracking and disturbance rejection, under sea-level and varying flight conditions, due to its global nature.
@@ 1 Introduction
[…]
@@
A plethora of GTE control literature is devoted to linear parameter-varying (LPV) models. To construct these models, linear time-invariant (LTI) subsystems are first identified around multiple steady-state points with subsequent interpolation or fitting of the LTI coefficients w.r.t. to some scheduling parameter, e.g. spool speed. For instance, Wei et al. [11] combined a piecewise linear model with nonlinear gains to construct an adaptive Wiener model. They results has shown that the Wiener approach can outperform a general describing function from [12] and the piecewise model without the nonlinear gains. The model is often converted to a polytopic structure suitable for design of robust controllers using quadratic programming and linear matrix inequalities as discussed in, e.g., [13, 14, 15, 16, 17], where [16] is a comprehensive review of GTE control strategies. In [17], the authors also introduce deep learning to engine control. This is a next broad class of nonlinear control approaches - the data-driven control.
@@
The need for component characteristics leads to the first major problem – the construction of some engines does not allow one to measure reliably all characteristics. For instance, the backward-flow combustion chambers makes it almost impossible to install temperature and other sensors, which are necessary for turbine measurements, without invasive operations such as drilling additional holes etc. which might compromise the structural integrity of the engine. Regarding the CFD estimation of compressor and turbine maps, these would require a detailed knowledge of the engine geometry, which is, in many cases, difficult to obtain from measurements or it is classified by the manufacturer. To overcome these problems, the identification of the GTE dynamical models solely from the measurable response data is often employed nowadays for the nonlinear control. Among the data-driven methods are, e.g., artificial neural networks (ANNs) [17, 18, 19] and their variation like recurrent networks, long short-term memory networks [20, 21], and auto-regressive nonlinear models with exogenous inputs [22, 23].
@@
Other methods focus on discovery of the underlying governing equations of the system from data. Sparse identification of nonlinear dynamics (SINDy) was developed by Brunton et al. at Washington University and first presented in [24]. The method looks for an interpretable governing equation approximated by a linear combination of nonlinear functions of states and inputs. Firstly, a library of candidate functions is designed, and the least squares (LS) problem is subsequently solved to obtain optimal weight coefficients. SINDy focuses on the trade-off between accuracy and interpretability of the model via the L1 regularization. The LS problem is usually solved using the LASSO algorithm [25]. SINDy was successfully used for identification of a small turbojet engine by authors in [26, 27], where the accuracy of the model was further enhanced using an EKF leading to a reliable prediction. Although SINDy is a powerful tool, the resulting model is still nonlinear. It can still be converted to control-affine form and utilized for MPC and design of EKF as mentioned above or feedback linearization, sliding mode control and more.
@@
The Koopman operator theory (KOT) represents a way to transform the nonlinear dynamics into a linear framework in an infinite-dimensional lifted state space [28]. In practice, a finite-rank truncation is demanded with a suitable set of observables. This allows for design of optimal controllers in the lifted state space using quadratic programming approaches known from the regulator optimization as demonstrated, e.g., in [29, 30, 31]. The lifting is performed using arbitrary nonlinear functions called observables, preserving the nonlinear nature. Although the definition of an optimal set of observables is still a problem-specific and open research question, several data-driven approaches exist to formulate a suitable set of observables without resorting to a trial-and-error approach.
Regarding the KOT, two ways can be distinguished as stated in the following list:
@@
The former includes methods such as extended DMD (EDMD) [32], exact DMD [33], generator EDMD [34], residual DMD [35], kernel DMD [36], Hankel DMD using Krylov sequences [28], and many more. The common goal is to find an optimal approximation of the Koopman operator in form of a finite-rank matrix via solution of a corresponding LS problem with additional constraints. The observables are constructed either by trial and error, building a large library of candidate observables, e.g. radial basis functions (RBFs) or polynomials, with a subsequent model-order reduction, or using some apriori knowledge about the system. Other methods seek to optimize the observables via machine learning [37, 38, 39, 40, 41]. For instance, the KOT with deep dictionary learning was utilized for identification and fault diagnostics of gas turbines in [41]. The observables were represented by ANN with an autoencoder architecture. A deep transfer learning framework was developed to adaptively capture the degraded states of the turbine. The method was key for extraction of a reliable nonlinear model. However, utilization of the autoencoder, despite being a powerful tool, leads to a complex model with a nonlinear inverse transformation back to the original state space making the control design still particularly challenging. Our previous work in [40] presented a way to optimize the observables using metaheuristic global optimization tools for the identification of GTEs. Results pointed to the suitability of the proposed method for discovery of low-order Koopman models for GTEs.
@@
The eigenfunction methods focus more on spectral properties of the Koopman operator. Under the assumption that the aforementioned matrix is diagonalizable, the underlying autonomous dynamics can be decomposed using a linear combination of spectral components represented by eigenvalues, eigenfunctions and eigenmodes. This was thoroughly discussed in literature and various papers and a book by Igor Mezić [42, 43, 44]. One of the rather rigorous spectral methods with a strong mathematical foundation is the generalized Laplace analysis [28]. The method is based on an iterative computation using time averages of a weighted signal obtained by subtraction of the previously computed modes from the measured trajectory. Korda et al. [45] proposed an optimal construction of Koopman eigenfunctions from data using a dense set of trajectories and mapping only the initial conditions to reveal the spatial structure of eigenfunctions. The method needs to be supplied by guessed eigenvalues which is the main drawback of this approach. The paper also presents ways to optimize the eigenvalue guesses, but it relies on the analytical gradient. These ideas were further extended also to the control synthesis in [31], where several methods were proposed including the implicit SINDy algorithm or a conversion to an equivalent eigenvalue problem for discovery of eigenfunctions. The method again assumed apriori knowledge of eigenvalues. The work showed that linear quadratic controllers can be effectively designed in eigenfunction coordinates using simple tools solving the state-dependent algebraic Riccati equation.
@@
  • The data required for the identification need to capture the autonomous behavior of the system. This is difficult to address for GTEs. A solution would be to accelerate the engine to the maximum operating point, cut off the fuel delivery, and measure the run-down curves. However, control systems of some engines include a cooling logic that keeps the rotor spinning at low RPM in order to actively cool hot sections. Deactivation of this logic could lead to a reduction of the remaining useful life of the engine.
@@
  • Some necessary performance data cannot be obtained due to the architecture of the engine not allowing for installation of sensors where needed. Thus, the CLM cannot be developed without additional procedures, such as matching of generic maps to meet the measured transient response of the GTE.
@@
  • Experimental derivation of LPV models without CLM is a time-consuming process leading only to a locally linearized dynamical model. Furthermore, the on-board adaptive nonlinear models, and the corresponding controllers, are often computationally expensive. Thus, it complicates their usage for small unmanned aerial vehicles where it might not be possible to install and maintain a processing unit with the required computational power.
@@
To address these challenges in this paper, the SINDy algorithm was first utilized to discover the underlying governing equation of the modeled single-spool GTE from data. The data can be measured during the normal operation of the engine and under the feedback control law with no need for adjustment of the existing control system. The testing signal should be designed such that the trajectory meets the control limits of the engine that are not known a priori, because the engine factory controller settings are not known.
@@
Since we do not currently have access to a measured engine response, the identification and analysis were conducted using an in-house nonlinear CLM of a generic single-spool GTE. This model was developed in the MATLAB® environment and validated against the GasTurb commercial engine simulation software [46]. A control-affine model was considered as a valid approximation of the GTE dynamics. Subsequently, the autonomous part of the SINDy model was utilized for identification of optimal spectral components yielding a conversion to the Koopman eigenfunction model (KEM). Genetic algorithm (GA) was employed to find optimal eigenvalues via the temporal projection of autonomous trajectories onto a span of exponential trajectories. Gradient-based algorithm was further developed to find the corresponding eigenfunctions. To enhance the prediction accuracy of the KEM, the Kalman observer was designed. Finally, an optimal Koopman Linear Quadratic Gaussian controller with Integral action (K-LQGI) was designed using global optimization methods and compared against a classical and gain scheduled LPV proportional-integral (PI) controllers.
@@
The rest of the paper is organized as follows: In Section 1, the introduction to engine modeling and control, and the problem formulation are provided. In Section 2, the reference physics-based model of the GTE is presented and briefly described. Sections 3 briefly introduces the Koopman operator. In sections 4 and 5, the main contributions of this paper, i.e. the SINDy-KEM-based identification and controller design, are delivered. Section 6 presents the main results of application of the proposed approach to a generic GTE. The KEM controller is compared to the classical PI and nonlinear LPV-PI controllers, and another innovation of this paper, the SINDy internal model controller, in both the sea-level and varying flight conditions. The final Section 7 summarizes the contributions and results.
@@ 2 Component-level model of a turbojet engine
To validate the proposed approaches, it was necessary to have control over the reference nonlinear model in order to connect various controllers and evaluate their performance. Thus, an in-house CLM of a generic engine was developed in the MATLAB R2022a environment. The CLM was further utilized for generation of the training and test datasets. It is crucial that the model is well-representative of a real engine and thus it was validated against the GasTurb 12 commercial simulation software, which represents the benchmark for engine modeling [46].
@@ 2.1 Models of GTE components
2.1.1 Atmospheric conditions
Atmospheric conditions were modeled in accordance with the international standard atmosphere. Concerning the pressure losses, these were considered in the inlet, combustion chamber (CC), exhaust duct and exhaust nozzle. The CC losses were modeled as a function of the CC inlet corrected mass flow parameter.
@@ 2.1.2 Compressor
Compressor model was represented by the compressor map, where the mass flow, pressure ratio and efficiency were converted to functions of the corrected spool speed and an auxiliary nonlinear coordinate - $\beta_{c}$ parameter. The detailed explanation of the modified $\beta_{c}$ interpolation approach is provided in [47]. Thus, the mass flow $W_{c}$, pressure ratio $\Pi_{c}$ and adiabatic efficiency $\eta_{c}$ were expressed as
$[W_{c,corr},\Pi_{c},\eta_{c}]=f(\beta_{c},N_{corr}),$ (1)
where $W_{c,corr}$ and $N_{corr}$ are the corrected compressor air mass flow and the corrected spool speed, respectively.
@@
where $C_{p,a}$ is the heat capacity of air and $\gamma_{a}$ is the adiabatic exponent of air.
@@ 2.1.3 Combustion chamber
Combustor model consisted of the map of burning efficiency $\eta_{b}$ as a function of the combustor loading parameter $\Omega_{cc}$, which was described, e.g., in [46, 48], and which accounts for the changes in the engine state and flight conditions. Regarding the pressure losses, these were modeled as a function of both the mass flow parameter and the mean temperature relative to the design point value. This way, the effect of heat addition on the decrease of the total pressure recovery was captured. The power balance of the combustion chamber is given as
$\eta_{b}W_{f}H_{u}=W_{cc}\bar{C}_{p}(T_{cc},T_{t})(T_{t}-T_{cc}),$ (3)
where $\eta_{b}$, $W_{f}$ and $H_{u}$ are the burning efficiency, fuel mass flow and lower heating value of fuel, respectively, and $\bar{C}_{p}$, $W_{cc}$, $T_{cc}$ and $T_{t}$ are the mean heat capacity of gases in CC, CC inlet air mass flow, CC inlet total temperature and turbine inlet total temperature, respectively.
This equation was utilized to determine the $T_{t}$ for a known $W_{f}$. The mean capacity $\bar{C}_{p}$ is a known function of both inlet and outlet conditions in CC, making the equation implicit. Thus, the fixed-point iteration was employed to solve for $T_{t}$. Since $\eta_{b}=f(\Omega_{cc})$, and $\Omega_{cc}=f(T_{cc},p_{cc},W_{cc})$, the efficiency can be rewritten as $\eta_{b}=f(T_{cc},p_{cc},W_{cc})$ and thus it is a known explicit function of CC inlet conditions.

### v6

# Koopman Eigenfunction-Based Identification and Optimal Nonlinear Control of Turbojet Engine

@@ Abstract
Gas turbine engines are complex and highly nonlinear dynamical systems. Deriving their physics-based models can be challenging because it requires performance characteristics that are not always available, often leading to many simplifying assumptions. This paper discusses the limitations of conventional experimental methods used to derive component-level and locally linear parameter-varying models, and addresses these issues by employing identification techniques based on data collected from standard engine operation under closed-loop control. The rotor dynamics are estimated using the sparse identification of nonlinear dynamics. Subsequently, the autonomous part of the dynamics is mapped into an optimally constructed Koopman eigenfunction space. This process involves eigenvalue optimization using metaheuristic algorithms and temporal projection, followed by gradient-based eigenfunction identification. The resulting Koopman model is validated against an in-house reference component-level model. A globally optimal nonlinear feedback controller and a Kalman estimator are then designed within the eigenfunction space and compared to traditional and gain-scheduled proportional-integral controllers, as well as a proposed internal model control approach. The eigenmode structure enables targeting individual modes during optimization, leading to improved performance tuning. Results demonstrate that the Koopman-based controller surpasses other benchmark controllers in both reference tracking and disturbance rejection under sea-level and varying flight conditions, due to its global nature.
@@ 1 Introduction
Gas turbine engines (GTEs) are highly nonlinear and control nonaffine systems, posing challenges for identification and control system design. The nonlinearities in GTE dynamics mainly result from aero-thermodynamic effects in turbomachinery components, ducts, volumes, combustion processes, and more [bib1]. A major category of GTE modeling and identification methods involves physics-based models. Among these, component-level models (CLMs) are popular, formed by connecting zero-dimensional models of engine components. These are usually represented by their performance characteristics, also known as maps, which are either measured on test stands or computed via computational fluid dynamics (CFD). These models can be used to derive linear parameter-varying (LPV) models for control design or employed directly as on-board estimators. The on-board real-time adaptive CLM was developed in [bib2] and further extended to limit protection [bib3], model predictive control (MPC) with extended Kalman filter (EKF) [bib4], and direct thrust control [bib5]. An on-board CLM with an unscented Kalman filter was studied in [bib6]. This combination demonstrated high estimation performance, and was further extended to the MPC framework and health parameter estimation [bib7]. A nonlinear dynamic-inversion controller with a dual EKF was proposed in [bib8], using a simplified physics-based model with approximate component maps. This model was later converted into a control-affine structure, which is broadly seen as a valid approximation of the actual engine dynamics. A sliding mode controller was used to improve the performance of a model-free adaptive controller for a variable-cycle GTE in [bib9]. A performance-seeking controller based on a dual-loop structure was introduced in [bib10]; the inner loop stabilizes the system, while the outer loop uses a CLM to optimize the GTE’s variable geometry, minimizing fuel consumption and addressing the multi-objective nature of GTE control. Results indicated that the controller reduced fuel consumption by 5 $\%$ while maintaining acceptable performance. On-board CLMs often require substantial computational resources, which may not be available on embedded hardware used on smaller flying platforms.
@@
A plethora of GTE control literature is devoted to linear parameter-varying (LPV) models. To construct these models, linear time-invariant (LTI) subsystems are first identified around multiple steady-state points with subsequent interpolation or fitting of the LTI coefficients w.r.t. to some scheduling parameter, e.g., spool speed. For instance, Wei et al. [bib11] combined a piecewise linear model with nonlinear gains to construct an adaptive Wiener model. The results have shown that the Wiener approach can outperform a general describing function from [bib12] and the piecewise model without the nonlinear gains. The model is often converted to a polytopic structure suitable for the design of robust controllers using quadratic programming and linear matrix inequalities as discussed in, e.g., [bib13, bib14, bib15, bib16, bib17], where [bib16] is a comprehensive review of GTE control strategies. In [bib17], the authors also introduce deep learning to engine control. This is the next broad class of nonlinear control approaches - the data-driven control.
@@
The need for component characteristics leads to the first major problem: the construction of some engines does not allow one to reliably measure all characteristics. For instance, backward-flow combustion chambers make it almost impossible to install temperature and other sensors, which are necessary for turbine measurements. The CFD estimation of compressor and turbine maps would require a detailed knowledge of the engine geometry which is either proprietary, or difficult to obtain from measurements. Additionally, it is challenging to estimate the possible dimension changes due to the thermal effects. To avoid these issues, the identification of GTEs rather relies on a black-box approach using dynamical response data that can be easily measured. Among the data-driven methods are, e.g., artificial neural networks (ANNs) [bib17, bib18, bib19] and their variants, like recurrent networks, long short-term memory networks [bib20, bib21], and auto-regressive nonlinear models with exogenous inputs [bib22, bib23].
@@
Other methods focus on the discovery of the underlying governing equations of the system from data. Sparse identification of nonlinear dynamics (SINDy) was developed by Brunton et al. at Washington University and first presented in [bib24]. The method looks for an interpretable governing equation approximated by a linear combination of nonlinear functions of states and inputs. Firstly, a library of candidate functions is designed, and the least squares (LS) problem is subsequently solved to obtain optimal weight coefficients. SINDy focuses on the trade-off between accuracy and interpretability of the model via the L1 regularization. The LS problem is usually solved using the LASSO algorithm [bib25]. SINDy was successfully used for the identification of a small turbojet engine by the authors in [bib26, bib27], where the accuracy of the model was further enhanced using an EKF, leading to a reliable prediction. Although SINDy is a powerful tool, the resulting model is nonlinear. Despite that, the model can still be converted to a control-affine structure and utilized for MPC, design of EKF as mentioned above, feedback linearization, sliding mode control, and more.
@@
The Koopman operator theory (KOT) represents a way to transform the nonlinear dynamics into a linear framework in an infinite-dimensional lifted state space [bib28]. In practice, a finite-rank truncation is demanded with a suitable set of transformation functions. This allows for the design of optimal controllers in the lifted state space using quadratic programming approaches known from the regulator optimization as demonstrated, e.g., in [bib29, bib30, bib31]. The lifting is performed using arbitrary nonlinear functions called observables, preserving the nonlinear nature. Although the definition of an optimal set of observables is still a problem-specific and open research question, several data-driven approaches exist to formulate a suitable set of observables without resorting to a trial-and-error approach.
Regarding the KOT, two approaches can be distinguished, as stated in the following list:
@@
The former includes methods such as extended DMD (EDMD) [bib32], exact DMD [bib33], generator EDMD [bib34], residual DMD [bib35], kernel DMD [bib36], Hankel DMD using Krylov sequences [bib28], and many more. The common goal is to find an optimal approximation of the Koopman operator in the form of a finite-rank matrix via solution of a corresponding LS problem with additional constraints. The observables are constructed either by trial and error, building a large library of candidate observables, e.g., radial basis functions (RBFs) or polynomials, with a subsequent model-order reduction, or using some a priori knowledge about the system. Some methods seek to optimize the observables via machine learning [bib37, bib38, bib39, bib40, bib41]. For instance, the KOT with deep dictionary learning was utilized for identification and fault diagnostics of gas turbines in [bib41]. The observables were represented by an ANN with an autoencoder architecture. A deep transfer learning framework was developed to adaptively capture the degraded states of the turbine. The method was key for the extraction of a reliable nonlinear model. However, utilization of the autoencoder, despite being a powerful tool, leads to a complex model with a nonlinear inverse transformation back to the original state space, making the control design particularly challenging. Our previous work in [bib40] presented a way to optimize the observables using metaheuristic global optimization tools for the identification of GTEs. Results pointed to the suitability of the proposed method for the discovery of low-order Koopman models for GTEs.
@@
The eigenfunction methods focus more on the spectral properties of the Koopman operator. Under the assumption that the aforementioned matrix is diagonalizable, the underlying autonomous dynamics can be decomposed using a linear combination of spectral components represented by eigenvalues, eigenfunctions, and eigenmodes. This was thoroughly discussed in literature, e.g., by Igor Mezić [bib42, bib43, bib44]. One of the rather rigorous spectral methods with a strong mathematical foundation is the generalized Laplace analysis [bib28]. The method is based on an iterative computation using time averages of a weighted signal obtained by subtraction of the previously computed modes from the measured trajectory. Korda et al. [bib45] proposed an optimal construction of Koopman eigenfunctions from data using a dense set of trajectories and mapping only the initial conditions to reveal the spatial structure of eigenfunctions. The method needs to be supplied with guessed eigenvalues, which is the main drawback of this approach. The paper also presents ways to optimize the eigenvalue guesses, but it relies on the analytical gradient. These ideas were also extended to the control synthesis in [bib31], where several methods were proposed, including the implicit SINDy algorithm or a conversion to an equivalent eigenvalue problem for the discovery of eigenfunctions. The method again assumed a priori knowledge of eigenvalues. The work showed that linear quadratic controllers can be effectively designed in eigenfunction coordinates using simple tools, solving the state-dependent algebraic Riccati equation.
@@
  • The data required for the identification should capture the autonomous behavior of the system. This is difficult to address for GTEs. A solution would be to accelerate the engine to the maximum operating point, cut off the fuel delivery, and measure the run-down curves. However, control systems of some engines include a cooling logic that keeps the rotor spinning at low RPM in order to actively cool hot sections. Deactivation of this logic could lead to a reduction of the remaining useful life of the specific engine.
@@
  • Some necessary performance data cannot be obtained due to the architecture of the engine not allowing for the installation of sensors where needed. Thus, the CLM cannot be developed without additional procedures, such as matching generic component maps to the measured transient response of the GTE.
@@
  • Experimental derivation of LPV models without CLM is a time-consuming process, leading only to a locally linearized dynamical model. Furthermore, the onboard adaptive nonlinear models and the corresponding controllers are often computationally expensive. This could complicate their usage, e.g. on small unmanned aerial vehicles, where it might not be possible to install and maintain a processing unit with the required computational power.
@@
To address these challenges in this paper, the SINDy algorithm was first utilized to discover the underlying governing equation of the modeled single-spool GTE from data. The data can be measured during normal engine operation under closed-loop control without modifying or disconnecting the existing control system. The testing signal should be designed in such a way that the controlled trajectory meets the engine operating limits.
@@
Since we do not currently have access to a measured engine response, the identification and analysis were conducted using an in-house nonlinear CLM of a generic single-spool GTE. This model was developed in the MATLAB® environment and validated against the GasTurb commercial engine simulation software [bib46]. A control-affine model was considered a valid approximation of the GTE dynamics. Subsequently, the autonomous part of the SINDy model was utilized for the identification of optimal spectral components, yielding a conversion to the Koopman eigenfunction model (KEM). A genetic algorithm (GA) was employed to find optimal eigenvalues using the temporal projection of autonomous trajectories onto a span of exponential trajectories. A gradient-based algorithm was further developed to find the corresponding eigenfunctions. To enhance the prediction accuracy of the KEM, a Kalman observer was designed. Finally, an optimal Koopman Linear Quadratic Gaussian controller with Integral action (K-LQGI) was designed using global optimization methods and compared against the classical LPV proportional and integral (PI) controller, and its gain-scheduled version.
@@
The rest of the paper is organized as follows: In Section 1, the introduction to engine modeling and control is provided, and the corresponding problems are formulated. In Section 2, the reference physics-based model of the GTE is presented and briefly described. Section 3 briefly introduces the Koopman operator. In Section 4 and Section 5, the main contributions of this paper, i.e., the SINDy-KEM-based identification and controller design, are presented. Section 6 presents the main results of the application of the proposed approach to a generic GTE. The KEM controller is compared to the classical PI and nonlinear LPV-PI controllers, and another innovation of this paper, the SINDy internal model controller, in both sea-level and varying flight conditions. The final Section 7 summarizes the contributions and results.
@@ 2 Component-level model of a turbojet engine
To validate the proposed approaches, it was necessary to have a full access to all parameters of the reference nonlinear model, and to have the option to connect various controllers for testing. Thus, an in-house CLM of a generic engine was developed in the MATLAB R2022a environment. The CLM was further utilized for the generation of the training and test datasets. To ensure that the model is a good representative of a real GTE, it was validated against the GasTurb 12 commercial simulation software, which is the benchmark for engine modeling [bib46].
@@ 2.1 Models of GTE components
2.1.1 Atmospheric conditions and losses
Atmospheric conditions were modeled according to the International Standard Atmosphere. The pressure losses were considered in the inlet, combustion chamber (CC), exhaust duct, and exhaust nozzle. The CC losses were modeled as a function of the CC inlet corrected mass flow parameter.
@@ 2.1.2 Compressor
Compressor model was represented by the compressor map, where the mass flow, pressure ratio, and efficiency were converted to functions of the corrected spool speed and an auxiliary nonlinear coordinate - $\beta_{c}$ parameter. The detailed explanation of the modified $\beta_{c}$ interpolation approach is provided in [bib47]. Thus, the mass flow $W_{\text{c}}$, pressure ratio $\mathrm{\Pi_{c}}$ and adiabatic efficiency $\eta_{c}$ were expressed as
$W_{c,corr},\Pi_{c},\eta_{c}]=f_{c}(\beta_{c},N_{corr}),$ (1)
where $W_{c,corr}$ and $N_{corr}$ are the corrected compressor air mass flow and the corrected spool speed, respectively. For more details about the GTE corrected quantities, the reader is referred to [bib1]. $N_{corr}$ will be defined later.
@@
where $W_{c}$ is now the physical air mass flow, $C_{p,a}$ is the heat capacity of air, and $\gamma_{a}$ is the adiabatic exponent of air.
@@ 2.1.3 Combustion chamber
The combustor model consisted of the map of burning efficiency $\eta_{b}$ as a function of the combustor loading parameter $\Omega_{cc}$ described, e.g., in [bib46, bib48], and which accounts for the changes in the engine state and flight conditions. Regarding the pressure losses, these were modeled as a function of both the mass flow parameter and the mean temperature relative to the design point value. This way, the effect of heat addition on the decrease of the total pressure recovery was captured. The power balance of the combustion chamber is given as
$\eta_{b}W_{f}H_{L}=W_{cc}\bar{C}_{p}(T_{cc},T_{t})(T_{t}-T_{cc}),$ (3)
where $\eta_{b}$, $W_{f}$ and $H_{L}$ are the burning efficiency, fuel mass flow and lower heating value of fuel, respectively, and $\bar{C}_{p}$, $W_{cc}$, $T_{cc}$ and $T_{t}$ are the mean heat capacity of gases in CC, CC inlet air mass flow, CC inlet total temperature and turbine inlet total temperature, respectively.
This equation was utilized to determine the $T_{t}$ for a known $W_{f}$. The mean capacity $\bar{C}_{p}$ is a known function of both inlet and outlet conditions in CC, making the equation implicit. Thus, the fixed-point iteration was employed to solve for $T_{t}$. Since $\eta_{b}=f_{b}(\Omega_{cc})$, and $\Omega_{cc}=f_{\Omega}(T_{cc},p_{cc},W_{cc})$, the efficiency can be rewritten as $\eta_{b}=f_{b,2}(T_{cc},p_{cc},W_{cc})$, making it a known explicit function of the CC inlet conditions.